"x^p+\frac{\frac{q}{2}y^{2p-q}}{(xy^2)^ppq}"


1
Expert's answer
2020-07-06T16:49:52-0400

"f(x,y)=x^p+\frac{\frac{q }{2}\times y^{2p−q}}{(xy^2)^p \times p q}" "(" Where "y" is a function of "x)"

"\frac{df}{d x}=px^{p-1}+" "\frac{(xy^2)^p\times pq(\frac{q}{2}(2p-q)y^{2p-q-1}\frac{dy}{dx})-\frac{q }{2}\times y^{2p−q}(pq(x^p(2p)y^{2p-1}\frac{dy}{dx}+px^{p-1}y^2p)) }{((xy^2)^p\times pq)^2}"